# Computation method for perfectoid purity and perfectoid BCM-regularity

### v4

@@ 2. Case(ii-ii):
If $\eta$ is in the case (ii-i), then $\Phi_{R,n}(\eta)=V^{n-1}(\eta_{1})$ for some $n\geq 1$. By the commutative diagram
@@
it is enough to show $V^{n-1}\colon H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{S})\to H^{d}_{\mathfrak{m}}(Q_{S,n})$ is injective. By the exact sequence in Proposition 4.3 and $H^{d-1}_{\mathfrak{m}}(\overline{S})=0$, we obtain $H^{d-1}_{\mathfrak{m}}(Q_{S,n-1})=0$. Therefore, we have $V^{n-1}\colon H^{d}_{\mathfrak{m}}(F^{n}_{*}\overline{S})\to H^{d}_{\mathfrak{m}}(Q_{S,n})$ is injective by Proposition 4.3. ∎
Proposition 5.7.
We use Notation 5.1. We assume $R$ is complete intersection. We take a surjection $S\to R$ and a perfectoid $R_{\infty}:=R^{S_{\infty}}_{perfd}$ as in [p-pure]*Lemma 4.23. We run a program in Definition 5.4 for an element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})\backslash\{0\}$ and assume that we obtain $\eta:=\eta_{0},\ldots,\eta_{i}$ for some $i\geq 0$ and $\eta_{i}$ is in the case of (i-ii). Then $\eta\notin\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))$.
@@
We set $\rho\colon R\to R_{\infty}$. We note that $H^{d-1}_{\mathfrak{m}}(\overline{R}_{\infty})=0$ by [p-pure]*Theorem 4.24. Suppose $\eta\in\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{\rho}))$, where $\overline{\rho}\colon\overline{R}\to\overline{R}_{\infty}$. Then $\eta_{i}\in\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{\rho}))$ by Lemma 5.6. Since we have $(\overline{R}_{\infty})_{perfd}\simeq\overline{R}_{perfd}$, there exists a positive integer $e^{\prime}$ such that $F^{e^{\prime}}(\eta_{i})=0$, it contradicts the fact that $\eta_{i}$ is in the case (i-ii). ∎
Theorem 5.8.
(cf. Theorem A) We use Notation 5.1. We assume $R$ is complete intersection. If $R$ is quasi-$(F,F^{\infty})$-split and $\operatorname{ht}(R)=n$, then $R$ is perfectoid pure and $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$, where $i_{n-1}:=1/p+\cdots+1/p^{n-1}$.
@@
By [p-pure]*Lemma 4.8 and Proposition 4.7, we may assume that $R$ is $\mathfrak{m}$-adic complete. We take a surjection $S\to R$ and a perfectoid $R_{\infty}:=R^{S_{\infty}}_{perfd}$ be as in [p-pure]*Lemma 4.23. Let $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ be a non-zero element of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$ and we run a program in Definition 5.4. If $\eta$ is in the case (i), then $\overline{R}$ is $F$-pure, and in particular, $\eta$ is in the case (i-ii). Thus, we obtain $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}((\overline{R}_{\infty})_{perf})$ is injective by the proof of Proposition 5.7. Since $(\overline{R}_{\infty})_{perf}\simeq R_{\infty}/(p^{1/p^{\infty}})$, $R$ is Gorenstein, and $H^{d}_{\mathfrak{m}}(R_{\infty})=0$, we have $(R,\varepsilon{\rm div}(p))$ is perfectoid pure for each $\varepsilon\in\mathbb{Z}[1/p]$ with $1>\varepsilon\geq 0$. In particular, we have $\operatorname{ppt}(R;{\rm div}(p))=1$. Next, we assume $\eta$ is in the case (ii), that is, $F(\eta)=0$, and in particular, we have $\operatorname{ht}(R)=n\geq 2$. Since $\eta$ is an element of the socle, we have $\Phi_{R,n}(\eta)\neq 0$ and $\Phi_{R,n-1}(\eta)=0$. Therefore, we obtain $\eta_{1}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $V^{n-1}(F^{n}_{*}\eta_{1})=\Phi_{R,n}(\eta)$. If there exists a positive integer $e$ such that $F^{e}(\eta_{1})=0$, then $\eta$ is in the kernel of the map
$\Phi_{R,n,e}\colon H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(Q_{R,n})\to H^{d}_{\mathfrak{m}}(Q_{R,n,e}),$
thus it contradicts the fact that $R$ is quasi-$(F,F^{\infty})$-split. Therefore, we obtain that $F^{e}(\eta_{1})\neq 0$ for every positive integer $e$, that is, $\eta_{1}$ is in the case of (i-ii), then by Lemma 5.6 and Proposition 5.7, we have $\eta$ is not contained in $\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))$, therefore by the above argument, $R$ is perfectoid pure. Next, let us compute $\operatorname{ppt}(R;{\rm div}(p))$. We take $\overline{\psi}_{n}\colon Q_{R_{\infty},n}\to\overline{R}_{\infty}$ and a unit $w\in\overline{R}_{\infty}$ as in Theorem 4.10. We consider the following commutative diagram;
[…]
where $\alpha$ follows from the natural surjection and the commutativity of right bottom diagram follows from Theorem 4.10. We consider the images of $\eta^{\prime}$ and $F^{-n}(\eta^{\prime}_{1})$ in $H^{d+1}_{\mathfrak{m}}(R_{\infty})$, then we obtain
(5.2) $p^{(1-i_{n})}(\eta^{\prime}/p)=wp^{i_{n-1}}F^{-n}(\eta^{\prime}_{1})/p^{i_{n}}=wF^{-n}(\eta^{\prime}_{1})/(p^{1/p^{n}}).$
Since $\eta^{\prime}_{1}$ is not contained in the kernel of $H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{1/p^{\infty}}))$ by $F^{e}(\eta)\neq 0$ for each $e\geq 1$, the element $F^{-n}(\eta^{\prime}_{1})$ is not contained in the kernel of $H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{1/p^{n}}))\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}/(p^{1/p^{\infty}})$, that is, we obtain $p^{\varepsilon}(F^{-n}(\eta^{\prime}_{1})/p^{1/p^{n}})\neq 0$ for every $1/p^{n}>\varepsilon\geq 0$. Therefore, by (5.2), we have $p^{\varepsilon}(\eta^{\prime}/p)=0$ if and only if $\varepsilon\geq 1-i_{n-1}$ for $\varepsilon\in\mathbb{Z}[\frac{1}{p}]$. In conclusion, we have $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$ by Proposition 2.9, as desired. ∎
@@
By Theorem 5.8, we can compute perfectoid pure thresholds for examples in Example 4.13 other than (1). We will compute perfectoid pure threshold of Example 4.13 (1) in Proposition 6.10.
@@
Theorem 5.10.
Proof.
We set $\overline{A}:=A/(p)$, $\overline{R}:=R/(p)$, $\overline{R}_{\infty}:=R_{\infty}/(p)$, $\overline{A}_{\infty}:=A_{\infty}/(p)$, and $d:=\dim\overline{R}$. We take a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. Suppose $R\to R_{\infty}$ is not perfectoid pure, then $\eta\in\operatorname{Ker}(H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty}))$. We run a program in Definition 5.4 for $\eta$ and get $\eta_{0}:=\eta,\eta_{1},\eta_{2},\ldots$ by Lemmas 5.6 and 5.7. We fix a system of $p$-power roots $\{f^{1/p^{e}}\}$ of $f$ in $A_{\infty}$. By the exact sequence
and $\tau_{i}^{(e)}:=\tau_{i}f^{1-1/p^{e}}$ for all integers $e,i\geq 0$. Let us show the following claim.
Claim 5.11.
We have $\tau_{i}^{(e)}\neq 0$ for all integers $e,i\geq 0$.
Proof.
Let us show the assertion by induction on $e$. Since $A$ is Cohen-Macaulay, the homomorphism $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d+1}_{\mathfrak{m}}(\overline{A})$ is injective. Since $A\to A_{\infty}$ is pure, the homomorphism $H^{d+1}_{\mathfrak{m}}(\overline{A})\to H^{d+1}_{\mathfrak{m}}(\overline{A}_{\infty})$ is injective, and in particular, we have $\tau_{i}^{(0)}=\tau_{i}\neq 0$ for every integer $i\geq 0$. Thus, we assume $e\geq 1$ and take $i\geq 0$. First, we assume that $\eta_{i}$ is in the case (i-i) in Definition 5.4. Then we have $F^{n}(\eta_{i})=\eta_{i+1}$ for some $n\geq 1$. By the commutative diagram
[…]
we have
$\tau_{i+1}^{(e-1)}$ $=u(\eta_{i+1})f^{1-1/p^{e-1}}=u\circ F^{n}(F^{n}_{*}\eta_{i})f^{1-1/p^{e-1}}$
$=F^{n}(F^{n}_{*}u(\eta_{i}))f^{p^{n}-1}f^{1-1/p^{e-1}}=F^{n}(F^{n}_{*}(u(\eta_{i})f^{1-1/p^{e-1+n}}))$
$=F^{n}(F^{n}_{*}\tau_{i}^{(e-1+n)}).$
By the induction hypothesis, we have $\tau_{i+1}^{(e-1)}\neq 0$, thus we have $\tau_{i}^{(e-1+n)}\neq 0$. Since $e-1+n\geq e$, we have $\tau_{i}^{(e)}\neq 0$, as desired. Next, we assume that $\tau_{i}$ is in the case (ii-i) in Definition 5.4. Then there exists an integer $n\geq 2$ such that $\Phi_{R,n}(\eta_{i})=V^{n-1}(F^{n}_{*}\eta_{i+1})$. We take regular sequence $x_{1},\ldots,x_{d}$ on $\overline{R}$ and elements $a,b\in\overline{R}$ such that $\eta_{i}=[a/x_{1}\cdots x_{d}]$, $\eta_{i+1}=[b/x_{1}^{p^{n}}\cdots x_{d}^{p^{n}}]$, and
$[a^{p}]\equiv V^{n-1}(b)\ \textrm{mod}\ ([x_{1}],\ldots,[x_{d}])Q_{R,n}.$
In particular, we have
By Claim 5.11, we have $\tau f^{1-p^{1/p^{e}}}=\tau_{0}^{(e)}\neq 0$ for every integer $e\geq 0$. Next, let us show $R\to R_{\infty}$ is pure. By the argument of [p-pure]*Proposition 6.5, it is enough to show that $fA\to(f^{1/p^{\infty}})$ is pure. Since $(f^{1/p^{\infty}})\simeq\mathrm{colim}_{e}f^{1/p^{e}}A_{\infty}$, it is enough to show that $fA\to f^{1/p^{e}}A_{\infty}$ is pure for every $e\geq 0$. We fix an integer $e\geq 0$. Since $f$ is a regular element in $A$ and $A_{\infty}$, it is enough to show that the homomorphism $A\to A_{\infty}\xrightarrow{\cdot f^{1-1/p^{e}}}A_{\infty}$ is pure. We consider the following commutative diagram in which horizontal sequence is exact:
[…]
Theorem 5.12.
Proof.
By Proposition 4.7 and [p-pure]*Lemma 4.13, we may assume that $R$ is complete. Therefore, there exist a regular local ring $R_{0}$ and regular sequence $f_{1},\ldots,f_{r}$ such that $R_{0}/(f_{1},\ldots,f_{r})\simeq R$. Let $R_{i}:=R_{0}/(f_{1},\ldots,f_{i})$ for $r\geq i\geq 1$. We take a perfectoid $R_{0,perfd}$ as in [p-pure]*Section 4.3. By Andre’s flatness lemma [BS]*Theorem 7.14, there exists a $p$-completely faithfully flat extension $R_{0,perfd}\to R_{0,\infty}$ to a perfectoid such that $R_{0,\infty}$ has systems of $p$-power roots of $f_{1},\ldots,f_{r}$. We set $R_{i,\infty}:=(R_{0,\infty}/(f_{1},\ldots,f_{i}))_{perfd}$. Let us show that $R_{i}\to R_{i,\infty}$ is pure for each $r\geq i\geq 0$ by induction on $i$. Since $R_{0}\to R_{0,\infty}$ is faithfully flat, we obtain the case of $i=0$. We take $i\geq 1$ and assume $R_{i-1}\to R_{i-1,\infty}$ is pure. By Theorem 5.10, it is enough to show that $R_{i-1,\infty}$ is cohomologically Cohen-Macaulay and $f_{i},p$ are regular elements of $R_{i-1,\infty}$. The cohomologically Cohen-Macaulayness follows from [p-pure]*Theorem 4.24. Furthermore, the sequence $p,f$ is a regular sequence on $R_{i-1,\infty}$ by the proof of [p-pure]*Theorem 4.24, and in particular, the element $p$ is a regular element. By Lemma 5.13, the element $f$ is a regular element of $R_{i-1,\infty}$, as desired. The last assertion follows from the first assertion and Proposition 4.6. ∎
Lemma 5.13.
Proof.
Lemma 5.14.
We use Notation 5.1. Let $R_{\infty}$ be a perfectoid over $R$ and $\eta\in H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})$. We assume $R_{\infty}$ has a system of $p$-power roots $\{p^{1/p^{e}}\}$ of $p$. If $\Phi_{R_{\infty},n}(\eta)\neq 0$ for some integer $n\geq 1$, then $p^{1-i_{n}}\eta\neq 0$, where $i_{n}=1/p+\cdots+1/p^{n}$.
Proof.
First, we consider the case where $n=1$. Then the image of $\eta$ by the composition
[…]
is non-zero. Indeed, the homomorphism $F^{-1}\colon\overline{R}_{\infty}\to R_{\infty}/(p^{1/p})$ is an isomorphism. Considering the images in $H^{d+1}_{\mathfrak{m}}(R_{\infty})$, we obtain $p^{1-1/p}(\eta/p)\neq 0$. In particular, we have $p^{1-1/p}\eta\neq 0$. Next, let us consider the case where $n\geq 2$. We may assume $\Phi_{R,n-1}(\eta)=0$. Then $p^{1-i_{n}}\eta\neq 0$ by the proof of Theorem 5.8, as desired. ∎
Theorem 5.15.
(cf. Theorems A and 6.9) We use Notation 5.1. We assume $R$ is complete intersection. Let $n$ a positive integer. Then $\operatorname{ht}(R)=n$ if and only if $R$ is perfectoid pure and
Proof.
We set $i_{m}:=1/p+\cdots+1/p^{m}$ for integers $m\geq 1$ and
$\varepsilon_{0}:=1-\frac{p+\cdots+p^{n-1}}{p^{n}-1}=1-i_{n}+1/p^{n}(1-i_{n})+1/p^{2n}(1-i_{n})+\cdots.$
Let us show the “only if” part. Thus, we assume $\operatorname{ht}(R)=n$. First, the inequality $\operatorname{ppt}(R;{\rm div}(p))\leq i_{n-1}$ follows from Proposition 5.2, thus let us show the other inequality. If $n=1$, then $\overline{R}$ is $F$-split, thus we have $\operatorname{ppt}(R;{\rm div}(p))=1$ by Theorem 5.8, as desired. We assume $n\geq 2$ and take a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. First, let us show that $\operatorname{ppt}(R;{\rm div}(p))\geq\varepsilon_{0}$. Then there exists $\eta_{1}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta)=V^{n-1}(F^{n}_{*}\eta_{1})$. We take a perfectoid $R_{\infty}$ over $R$ as in the proof of Theorem 5.8. The image of $\eta_{1}$ by $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})$ is denoted by $\eta^{\prime}_{1}$. If $p^{\alpha}\eta^{\prime}_{1}\neq 0$ for some $\alpha>\varepsilon_{0}$, then we have
$\operatorname{ppt}(R;{\rm div}(p))\geq 1-i_{n}+\alpha/p^{n}>\varepsilon_{0}$
by the proof of Theorem 5.8. Therefore, we may assume $\Phi_{R_{\infty},n-1}(\eta^{\prime}_{1})=0$ by Lemma 5.14. Since $R$ is perfectoid pure by Theorem 5.12, we obtain $\Phi_{R,n-1}(\eta_{1})=0$. Therefore, we obtain $\eta_{2}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta_{1})=V^{n-1}(F^{n}_{*}\eta_{2})$. Repeating such a process, we may assume that we obtain a sequence $\eta_{0}:=\eta,\eta_{1},\eta_{2},\ldots$ on $H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta_{i})=V^{n-1}(F^{n}_{*}\eta_{i+1})$ for every $i\geq 0$. We set
where $\eta^{\prime}_{i}$ is the image of $\eta_{i}$ by $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})$ for every $i\geq 0$. By the proof of Theorem 5.8, we obtain $\alpha_{i}\geq 1-i_{n}$ and $1-i_{n}+\alpha_{i+1}/p^{n}=\alpha_{i}$ for every $i\geq 0$. In other words, we obtain $\alpha_{i+1}-\varepsilon_{0}=p^{n}(\alpha_{i}-\varepsilon_{0})$. Suppose $\alpha_{0}<\varepsilon_{0}$, then for enough large $i$, we have $\alpha_{i}<1-i_{n}$, thus we have a contradiction. Therefore, we obtain $\alpha_{0}=\operatorname{ppt}(R;{\rm div}(p))\geq\varepsilon_{0}$, as desired.
Next, let us show the “if” part. Since we have
$\operatorname{ppt}(R;{\rm div}(p))\geq\varepsilon_{0}>1-i_{n},$
we obtain $\operatorname{ht}(R)\leq n$ by Proposition 5.2. Furthermore, if $n\geq 2$ and $\operatorname{ht}(R)\leq n-1$, then we obtain a contradiction by the “only if” part. Therefore, we obtain $\operatorname{ht}(R)=n$, as desired. ∎
Corollary 5.16.
(cf. Theorem A) We use Notation 5.1. We assume $R$ is complete intersection. Let $n$ a positive integer. Then $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$ if and only if $R$ is quasi-$(F,F^{\infty})$-split of height $n$.
Proof.
The “if” part follows from Theorem 5.8. We assume $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$, then $\operatorname{ht}(R)=n$ by Theorem 5.15. We take a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. Then there exists $\eta_{1}\in H^{d}_{\mathfrak{m}}(\overline{R})$ such that $\Phi_{R,n}(\eta)=V^{n-1}(F^{n}_{*}\eta_{1})$. We take a perfectoid $R_{\infty}$ over $R$ as in the proof of Theorem 5.8. Since $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$, we have $p^{\alpha}\eta_{1}^{\prime}\neq 0$ for $1>\alpha\geq 0$ by the proof of Theorem 5.8, where $\eta^{\prime}_{1}$ is the image of $\eta_{1}$ by $H^{d}_{\mathfrak{m}}(\overline{R})\to H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})$. For an integer $e\geq 1$, since the composition
$H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})\xrightarrow{F^{e}}H^{d}_{\mathfrak{m}}(F^{e}_{*}\overline{R}_{\infty})\xrightarrow{F^{-e}}H^{d}_{\mathfrak{m}}(R_{\infty}/(p^{1/p^{e}}))$
coincides with the natural homomorphism $H^{d}_{\mathfrak{m}}(\overline{R}_{\infty})\to H^{d}_{\mathfrak{m}}(R_{\infty}/(p^{1/p^{e}}))$ and $F^{-e}\colon\overline{R}_{\infty}\to R_{\infty}/(p^{1/p^{e}})$ is an isomorphism, we obtain $F^{e}(\eta_{1}^{\prime})\neq 0$. Therefore, we obtain $\Phi_{R,n,e}(\eta)\neq 0$ for every $e\geq 1$, thus $R$ is quasi-$(F,F^{\infty})$-split of height $n$, as desired. ∎

### v5

@@ 5.2. Computation of perfectoid threshold via quasi-$F$-splitting
where $\alpha$ follows from the natural surjection and the commutativity of right bottom diagram follows from Theorem 4.10. We consider the images of $\eta^{\prime}_{j}$ and $F^{-n}(\eta^{\prime}_{j+1})$ in $H^{d+1}_{\mathfrak{m}}(R_{\infty})$, then we obtain
(5.2) $p^{(1-i_{n})}(\eta^{\prime}_{j}/p)=wp^{i_{n-1}}F^{-n}(\eta^{\prime}_{j+1})/p^{i_{n}}=wF^{-n}(\eta^{\prime}_{j+1})/(p^{1/p^{n}}).$
We consider the diagram
@@
where $\mathfrak{m}^{\flat}$ is the ideal of $R_{\infty}^{\flat}$ containing $T$ and $\mathfrak{m}^{\flat}/(T)\simeq\mathfrak{m}/(p)$, then the corresponding element to $\eta_{j}^{\prime}/p$ and $\eta_{j+1}^{\prime}/p$ are denoted by $\tau_{j}^{\flat},\tau_{j+1}^{\flat}\in H^{d+1}_{\mathfrak{m}}(R_{\infty}^{\flat})$, respectively. By (5.2), we have
$T^{(1-i_{n})}\tau_{j}^{\flat}=w^{\prime}F^{-n}(\tau_{j+1}^{\flat})$
for some unit $w^{\prime}\in R_{\infty}^{\flat}$. Since $F^{-n}$ is an isomorphism, we have $T^{\alpha}\tau_{j+1}^{\flat}=0$ if and only if $T^{1-i_{n}+\alpha/p^{n}}\tau^{\flat}_{j}=0$ for $\alpha\in\mathbb{Z}[1/p]_{\geq 0}$. In particular, we have $p^{\alpha}\eta^{\prime}_{j+1}=0$ if and only if $p^{1-i_{n}+\alpha/p^{n}}\eta_{j}^{\prime}=0$. Therefore, we have
$\alpha_{j}=1-i_{n}+p^{-n}\alpha_{j+1},$
as desired. $\blacksquare$
We write the $p$-adic expansion of $\alpha_{j}$ as
$\alpha_{j}=\sum_{m\geq 1}\frac{a_{j,m}}{p^{m}},$
where $a_{j,m}\in\{0,1,\ldots,p-1\}$ and the sum is not finite. We note that since $1\geq\alpha_{j}$, we can take such an expansion. Since we have
$1-i_{n_{j}}=\frac{p-2}{p}+\cdots+\frac{p-2}{p^{n_{j}-1}}+\frac{p-1}{p^{n_{j}}},$
we obtain
$a_{jm}=\begin{cases}p-2&\text{if $m\leq n_{j}-1$}\\ p-1&\text{if $m=n_{j}$}\\ a_{j+1,m}&\text{if $m\geq n_{j}+1$}\end{cases}$
by Claim 5.10. Therefore, we have
$a_{0,m}:=\begin{cases}p-1&\text{if $m=\sum_{i=0}^{r}n_{i}$ for some $r\geq 0$}\\ p-2&\text{otherwise},\end{cases}$
as desired. ∎
Theorem 5.11.
(cf. Theorems A and 6.9) We use the notation introduced in Notation 5.1 We assume $R$ is complete intersection. Let $n$ a positive integer. Then $\operatorname{ht}(R)=n$ if and only if $R$ is perfectoid pure and
@@
We take a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. To prove the “only if” direction, we assume that $\operatorname{ht}(R)=n$. We construct a sequence $\{\eta_{i}\}_{i\geq 0}$ with $\eta_{0}=\eta$ and $\{n_{i}\}_{i\geq 0}$ using the procedure in Definition 5.4. Then we have $n_{i}\leq n$ for every $i\geq 0$ and $n_{0}=n$. By Theorem 5.9, we have
$\sum_{m\geq 1}\frac{p-2}{p^{m}}+\sum_{m\geq n}\frac{1}{p^{m}}\geq\operatorname{ppt}(R;{\rm div}(p))\geq\sum_{m\geq 1}\frac{p-2}{p^{m}}+\sum_{r\geq 1}\frac{1}{p^{rn}}.$
By computation, we have
$1-\frac{1}{p}-\cdots-\frac{1}{p^{n-1}}\geq\operatorname{ppt}(R;{\rm div}(p))\geq 1-\frac{p+\cdots+p^{n-1}}{p^{n}-1},$
as desired.
Next, we prove the “if part”. Since we have
$\operatorname{ppt}(R;{\rm div}(p))\geq 1-\frac{p+\cdots+p^{n-1}}{p^{n}-1}>1-i_{n},$
we obtain $\operatorname{ht}(R)\leq n$ by Proposition 5.2. We construct a sequence $\{\eta_{i}\}_{i\geq 0}$ with $\eta_{0}=\eta$ and $\{n_{i}\}_{i\geq 0}$ using the procedure in Definition 5.4. Therefore, we have $n_{0}=n$ by Theorem 5.9, and in particular, we obtain $\Phi_{R,n}(\eta)\neq 0$ and $\Phi_{R,n-1}(\eta)\neq 0$ if $n\geq 2$. Since $\eta$ is an element of the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$, we have $\operatorname{ht}(R)=n$, as desired. ∎
Corollary 5.12.
(cf. Theorem A) We use the notation introduced in Notation 5.1 We assume $R$ is complete intersection. Let $n$ a positive integer. Then $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$ if and only if $R$ is quasi-$(F,F^{\infty})$-split of height $n$, where $i_{n-1}=1/p+\cdots+1/p^{n-1}$.
@@
We take a non-zero element $\eta\in H^{d}_{\mathfrak{m}}(\overline{R})$ in the socle of $H^{d}_{\mathfrak{m}}(\overline{R})$. To prove the “if” part, we assume that $R$ is quasi-$(F,F^{\infty})$-split of height $n$. We construct a sequence $\{\eta_{i}\}_{i\geq 0}$ with $\eta_{0}=\eta$ and $\{n_{i}\}_{i\geq 0}$ using the procedure in Definition 5.4. By the definition of quasi-$(F,F^{\infty})$-split, we have $n_{0}=n$ and $n_{i}=1$ for every $i\geq 1$. Thus, we obtain $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$ by Theorem 5.9, as desired.
To prove the “only if’ part, we assume $\operatorname{ppt}(R;{\rm div}(p))=1-i_{n-1}$, then $\operatorname{ht}(R)=n$ by Theorem 5.11. We construct a sequence $\{\eta_{i}\}_{i\geq 0}$ with $\eta_{0}=\eta$ and $\{n_{i}\}_{i\geq 0}$ using the procedure in Definition 5.4. By Theorem 5.9, we have $n_{0}=n$ and $n_{i}=1$ for $i\geq 1$. Therefore, the ring $R$ is quasi-$(F,F^{\infty})$-split of height $n$, as desired. ∎
@@
By Corollary 5.12, we can compute perfectoid pure thresholds for examples in Example 4.13 other than (1). We will compute perfectoid pure threshold of Example 4.13 (1) in Proposition 6.10.
@@
Corollary 5.15.
(Theorem B) We use the notation introduced in Notation 5.1 We assume $R$ is complete intersection and perfectoid pure. If $\operatorname{ppt}(R;{\rm div}(p))>\frac{p-2}{p-1}$, then there exists a positive integer $n$ such that $R$ is quasi-$n$-split, and in particular, we have
Proof.
Since we have
$\operatorname{ppt}(R;{\rm div}(p))>\frac{p-2}{p-1}=\lim_{n\to\infty}(1-i_{n}),$
there exists a positive integer $m$ such that $\operatorname{ppt}(R;{\rm div}(p))>1-i_{m}$. By Proposition 5.2, we have that $R$ is $m$-quasi-$F$-split, and in particular, there exists a positive integer $n$ such that $\operatorname{ht}(R)=n$. By Theorem A, we have
$1-\frac{1}{p}-\cdots-\frac{1}{p^{n-1}}\geq\operatorname{ppt}(R;{\rm div}(p))\geq 1-\frac{p+\cdots+p^{n-1}}{p^{n}-1},$
as desired. ∎